OpenStudy (anonymous):

What is the equation of the line that passes through the point (3, 2) and has a slope of 4?

OpenStudy (anonymous):

The slope is 4 so the equation is : \[y=4x+b\] To find b, we have that the point (3,2) is from the line , so : \[2=4\times 3+b\] So : \[b=2-4\times3=-10\] So the equation is : \[y=4x-10\]

OpenStudy (anonymous):

The slope-intercept form is y = mx + b m is the slope b is the y-intercept form y = 4x + b 2 = 4(3) + b 2 = 12 + b b = 2-12 = -10 y = 4x - 10

OpenStudy (anonymous):

thank you for the help Guys!!!